# Investigating the effects of QCD matter’s electrical conductivity on charge dependent directed flow

### v1

@@ I Introduction
In the next section, Sec. II, we introduce our relativistic resistive magneto-hydrodynamic (RRMHD) model. Then in Sec. III we will show the time evolution of the magnetic field in our model. Next, how we fix the initial conditions of the EM fields as solutions of Maxwell’s equations in a medium and the initial conditions of QGP are discussed in Sec. IV. Finally, in Sec. VI we present the results for charge dependent directed flow using our RRMHD model. Our conclusions are summarized in Sec. VII and we discuss the necessary future improvements to move this type of study from qualitative to quantitative.
@@ II Hydrodynamic formalism of quark-gluon plasma with conductivity
For this work, we have built upon the relativistic resistive magneto-hydrodynamic (RRMHD) model of [21, 23, 22] such that we can study central collisions. The overall algorithms of the model remain the same, but some implementations have been changed to avoid sources of numerical instability; unrealistic negative pressures and imaginary Lorentz factors. We have also reduced the computation time by changing the numerical implementation to use the parallel computation architecture called Message Passing Interface (MPI). Because of the implementation changes, we have rebenchmarked the model using the same tests as described in the original paper and found only small differences. That model simultaneously solves the relativistic ideal-hydrodynamic equations and Maxwell’s equations with Ohm’s law as a current source. Reviewing the relevant details, we start with the conservation laws of ideal-hydrodynamics:
@@
$J^{\mu}=qu^{\mu}+\sigma F^{\mu\nu}u_{\nu}.$ (5)
We have assumed the electric conductivity $\sigma$ is a constant and $q=-J^{\mu}u_{\mu}$ is the electric charge density in the fluid’s comoving frame.
The feature of RRMHD compared to ideal relativistic magneto-hydrodynamics (MHD) is a finite conductivity $\sigma=1/\eta$, which is the inverse of $\eta$ the resistivity. In ideal relativistic MHD models of QGP, the conductivity is taken to be infinite, in other words, QGP is treated as a perfect conductor. However, Lattice QCD and pQCD calculations suggest the conductivity should be $\sigma=\mathcal{O}(0.1\sim 10)$ for the highest energy regions produced at RHIC and the LHC [7]. Additionally, a finite conductivity reduces the lifetime of the EM fields compared to idea MHD and can introduce new phenomena like magnetic reconnection. Modeling QGP with a finite conductivity leads to interesting physics beyond the ideal MHD scenario.
@@ III Time evolution of the magnetic field
The magnetic field plays a major role in the charge-dependent directed flow $\Delta v_{1}$ at symmetric collisions. One role the magnetic field plays is to induce Faraday and Hall effect currents in the QGP. Then the interplay of Lorentz + Coulomb + Faraday currents within the plasma leads to an electric chemical potential. During the time evolution, the chemical potential is frozen into the freezeout hypersurface. Thus, an understanding of the time evolution of the magnetic field is important when calculating an observable like $\Delta v_{1}$ which depends on freezeout hypersurface.
@@
  • We assume the electric conductivity of QGP is a constant value over all spacetime. Strictly speaking, dimensional analysis alone suggests the electric conductivity is temperature-dependent, which means it should change as the plasma expands and cools.
  • We ignore the pre-equilibrium dynamics that can create an initial electric charge distribution and change the configuration of the initial EM fields [26, 27].
  • Last, the initial EM fields are created by the spectators of the collision, but during the time evolution we only include the plasma charges as a source of the EM fields. In other words, our model focuses on the behavior of charges inside the plasma and how they alone affect the time evolution of the EM fields.
For a more complete description, we plan to resolve the first and third items in a future version of our model. The second item depends on an understanding of the pre-equilibrium dynamics, which has recently been an active area of research (e.g.,  [28, 29]). Some possible directions include Color Class Condensate (CGC), IP-Glasma, flux tube, and kinetic models.
@@
  • For, $\tau^{-1}+\sigma>1$ the magnetic field is over damped and exponentially decays.
  • At, $\tau^{-1}+\sigma=1$ the magnetic field is critically damped and will reach an asymptotic value as quickly as possible.
  • Then for, $\tau^{-1}+\sigma<1$ the magnetic field is under damped and will oscillate about an asymptotic value.
[…]
@@ IV Numerical setup and methods
For relativistic heavy-ion collisions, the symmetry of the system is best captured with Milne coordinates $(\tau,x,y,\eta)$, which have the metric $g_{\mu\nu}=\text{diag}(-1,1,1,\tau^{2})$. Accordingly, our RRMHD model solves Eqs.(1)-(5) using those coordinates with the volume factor $\sqrt{-g}=\tau$. We have adopted the optical Glauber model for the initial energy density profile and included an initial longitudinal tilt [31]. Even though we initialize the model after QGP is created at $\tau_{0}=0.4$ fm/$c$, the source of the initial EM fields is dominated by the collisions spectators [24]. Therefore, we calculate the initial EM fields using the analytic solution from [25], which is a solution of Maxwell’s equations in a medium with the ion charges as the source. These are the same initial conditions as the previous works [23, 21, 20] and are summarized in table 1.
@@
Numerically, we use a constant grid size of $168\times 168\times 72$ with 16.8 $-16.8\text{ fm}<(x,y)<16.8\text{ fm}$ and $-7.2<\eta<7.2$, except for calculating the initial EM fields. Instead, for the initial EM fields, we adopt the following approach based on the analytic solution [25].
@@
    Assume the nuclei are very thin in the beam direction $\eta$ because of Lorentz contraction, such that their charge is distributed only in the transverse plane $x,y$. Then we can fix the $\eta$ position of the charges based on the collision kinematics.
@@
    At each grid point $\begin{pmatrix}x,&y,&\eta\end{pmatrix}$, calculate the EM fields by integrating over the contribution from all possible sources. Practically, this means we create a new grid for the EM sources in the transverse plane of size $200\times 200$ for 15.0 $-15.0\text{ fm}<(x,y)<15.0\text{ fm}$ and integrate over it with the trapezoidal method. This size ensures the area of both nuclei are integrated over.
The second step is repeated for all the original grid points $\begin{pmatrix}x,&y,&\eta\end{pmatrix}$. The numerical Courant-Friedrichs-Lewy (CFL) value is fixed to be $C_{\text{CFL}}=0.05$.
Lastly, we calculate the final state hadron distribution using the Cooper-Frye formula [33]. We are ignoring the final state interactions that would be included in an afterburner, because our current goal is to focus on the effects of the QGP conductivity. Our implementation of the Cooper-Frye formula starts by constructing a 4-D hypersurface at the energy density boundary $\epsilon_{f}=0.15\text{ GeV/fm}^{3}$ [34]. We then construct a distribution using the fluid velocity $v_{i}$, and charge density $q$ that is perpendicular to the hypersurface $d\Sigma(t,x,y,z)$. Next, the accumulated charge density $q$ is converted to the electric chemical potential using the linear approximation $\mu_{Q}\approx q(\hbar c)^{2.5}/T_{f}^{2}$. Where the final temperature is calculated using the ideal gas equation of state,
$T_{f}=\left(\frac{\epsilon_{f}(\hbar c)^{3}}{3}\right)\approx 140.0\text{ MeV.}$. (18)
Then we integrate over the entire hypersurface to arrive at a distribution. The details of the implementation are given in appendix B.
@@ V RRMHD Results: Time evolution of the magnetic field
As discussed at the beginning of Sec. III, the time evolution of the magnetic field induces Lorentz + Faraday currents in QGP. Coulomb currents caused by the electric field of the positive spectators also appear, but for symmetric collisions largely cancel each other. Then the EM field component in the charge dependent directed flow $\Delta v_{1}$ for symmetric collisions is dominated by the Lorentz + Faraday currents, specifically the magnetic field. In particular, the $B_{y}$ component of the field has the largest contribution to $\Delta v_{1}$, which approximately appears as $\gamma B_{y}$ in the Faraday current and $-\gamma u_{z}B_{y}$ in the Lorentz current [15]. Where $\gamma$ is the Lorentz factor of QGP in the local fluid rest frame and $u_{z}$ is the $z$ component of the fluid velocity.
In our RRMHD model, the time evolution of $B_{y}$ depends on the initial conditions and the electric conductivity $\sigma$ of QGP. As a starting point, we have used smooth initial solutions to Maxwell’s equations in a medium as discussed in Sec. IV that predicts a large initial field intensity for $B_{y}$. That being said, solving the time evolution of the magnetohydrodynamics Eqs. (1)-(5) with different $\sigma$’s has an impact on $\Delta v_{1}$ through changes of $B_{y}$, which will be shown in Sec. VI.
@@
Figure 1: Time evolution of the $y$-component of the magnetic field at the center of the grid using our RRMHD model. We have assumed the electric conductivity of the plasma $\sigma$ is a constant in spacetime, and the spectators of the collision only act as an initial source for the field. The magnetic field’s time evolution is similar to a non-linearly damped oscillator. For conductivity values between, $10<\sigma<$ 1 the magnetic field starts under damped and oscillates around zero. Larger values of $\sigma$ cause the magnetic field to be over damped, which means it never crosses zero.
The reduction in differences between the results for larger sigma values, like between $\sigma=10\text{ fm}^{-1}$ and $\sigma=100\text{ fm}^{-1}$, is partly due to numerical diffusion overwhelming the physical diffusion. This is a long-standing issue for resistive (relativistic) MHD models in astrophysics, where the conductivities can become relativity large, see Ref. [35] as an example of a detailed analysis. Here for each simulation, we keep the typical length-scale $L$ of the problem fixed and the order of the fluid velocities $u$ remains largely unchanged, but the conductivity increases by an order of magnitude. This means the magnetic Reynolds number $S\equiv\sigma L|u|$ also increases by an order of magnitude for each conductivity. When studying QGP, we limit the conductivity to be $\sigma<5\text{ fm}^{-1}$. This limit is justified by Lattice, pQCD, and AdS/CFT calculations and should keep the numerical diffusion smaller than the physical [4, 5, 6, 7, 8, 9].
A useful feature of the analytic solutions used for the initial EM fields, see Sec. IV and Ref. [25], is that it can be applied to all times. As a comparison in Fig. 2 that analytic solution for $-B_{y}$ follows a similar time evolution as our RRMHD solution. The deviations after $t\sim 1.25$ fm/$c$ between the analytic and RRMHD solutions can be attributed to different EM sources used in each calculation. At early times, both our RRMHD model and the analytic calculation only use the spectators as sources of the EM fields. Then, during the time evolution, our RRMHD model only considers QGP as a possible EM source; whereas, the analytic calculation only includes the spectators.
@@ VI RRMHD Results: Charge dependent directed flow
[…]
which is dependent on $Y$ the momentum-space rapidity. As explained in Sec. II and appendix B we have adopted the Cooper-Frye formula [33, 34] for hadronization at the linear order approximation for the electric chemical potential $\mu_{Q}\simeq q/(g_{h}T_{f}^{2})$. Where $q$ is the electric charge density and $T_{f}=140$ MeV is the freezeout temperature. Finally, we use the Gauss-Legendre method to integrate over the transverse momentum $p_{T}$ and angle $\phi$ in Eq.(20).
@@
Figure 3: Electric chemical potential $\mu_{Q}$ that is accumulated on the freezeout hypersurface for $T_{f}=0.140$ GeV, $-0.2<\eta<0.0$, and $\tau>0.4$ fm/$c$. Outer points correspond to an earlier freezeout time, and the evolution steps toward the center of the grid. Because we start from smooth initial conditions, the transverse plane has reflection symmetry about zero on the $x$-axis and $y$-axis. The difference in magnitude of $\mu_{Q}$ between the panels is attributed to the value of QGPs electric conductivity $\sigma$.
For our model, the source of non-vanishing $\Delta v_{1}$ is a net electric charge $q$ on the freezeout hypersurface, which depends on the conductivity $\sigma$ through Eq.(5). As we have already discussed, the net electric charge is produced by the Lorentz and Faraday currents, and the initial positive charge of the plasma. Figures 3 and 4 illustrate the net electric chemical potential on the freezeout hypersurface. For a larger conductivity, the magnitudes of $\mu_{Q}$ increase because the lifetime of the EM fields have also increased, see Fig. 1 and arguments about the Lorentz and Faraday currents.
Figure 4: Longitudinal dependence of the electric chemical potential $\mu_{Q}$ that is accumulated on the freezeout hypersurface at $T_{f}=0.140$ GeV, -0.2 fm $<y<$ 0.0 fm, and $\tau>0.4$ fm/$c$. Outer points correspond to an earlier freezeout time, and the evolution steps toward the center of the grid. For peripheral collisions, the magnetic fields in the longitudinal plane $\eta$ are asymmetric about zero [22]. That leads to an asymmetry in $\mu_{Q}$ across $\eta$, which appears on the freezeout hypersurface. A 180 degrees rotation symmetry about the center $(0,0)$ exists because we start from smooth initial conditions.
@@
Additionally, an asymmetry in the chemical potential $\mu_{Q}$ is present in the longitudinal plane projected on Fig. 4. This asymmetry is created in peripheral collisions, i.e., $b\neq 0$ fm, because the initial magnetic fields of the spectators break the reflection symmetry across $\eta=0$. Consequently, the charged directed flow of Eq.(20) is different for positively vs. negatively charged final state hadrons.
@@
An observable difference between the charge of protons and anti-protons is illustrated in Fig. 5, where the conductivity of QGP is varied across the panels. The rapidity $Y$ dependence of (anti-)proton charge dependent directed flow Eq.(20) separates in panels b and c. That separation is captured by a nonzero difference $\Delta v_{1}\neq 0$ from Eq.(19). This is a similar result as the observations at the STAR experiment [16].
@@
The negative linear slope of $\Delta v_{1}$ in panels b and c of Fig. 5 is a sign that the Faraday current has a greater net effect than the Lorentz current, which is in agreement with Ref. [15]. This shows the approximation of the Faraday current as $\gamma B_{y}$ and the Lorentz current as $-\gamma u_{z}B_{y}$ at the beginning of Sec. III holds for our RRMHD model. Moreover, an order increase in the electric conductivity $\sigma$ brought a similar increase in the slope, $\Delta(dv_{1}/dy)$ suggesting this observable could be used to understand $\sigma$, the electric conductivity of QGP. Although, a more refined model is required to draw any quantitative conclusions.

### v2

# Investigating effects of the electrical conductivity of QCD matter on charge-dependent directed flow

@@ I Introduction
In the next section, Sec. II, we introduce our relativistic resistive magneto-hydrodynamic (RRMHD) model. Then in Sec. III we will discuss the time evolution of the magnetic field in our model. Next, how we fix the initial conditions of the EM fields as solutions of Maxwell’s equations in a medium and the initial conditions of QGP are discussed in Sec. IV. Results for the time evolution of the magnetic field are shown in Sec. V, and we discuss how the results are related to the initial conditions. In Sec. VI we present the results for charge dependent directed flow using our RRMHD model, including details on the connection with charge accumulation on the freezeout hypersurface. Our conclusions are summarized in Sec. VII and we discuss the necessary future improvements to move this type of study from qualitative to quantitative.
@@ II Hydrodynamic formalism of quark-gluon plasma with conductivity
For this work, we have built upon the relativistic resistive magneto-hydrodynamic (RRMHD) model  [21, 23, 22] such that we can study central collisions. That model simultaneously solves the relativistic ideal-hydrodynamic equations and Maxwell’s equations with Ohm’s law as a current source. Reviewing the relevant details, we start with the conservation laws of ideal-hydrodynamics:
@@
$J^{\mu}$ $=qu^{\mu}+\sigma F^{\mu\nu}u_{\nu},$ (5)
$=\gamma\begin{pmatrix}q+\sigma v_{j}E^{j},&qv^{i}+\sigma[E^{i}+\epsilon^{ijk}v_{j}B_{k}]\end{pmatrix}$, (6)
and use the ideal equation of state (EoS). We have assumed the electric conductivity $\sigma$ is a constant and $q=-J^{\mu}u_{\mu}$ is the electric charge density in the fluid’s comoving frame. Following the naming convention of Ref. [15], the $E^{i}$ of Eq.(6) is the Coulomb + Faraday currents and $\epsilon^{ijk}v_{j}B_{k}$ is the Lorentz current.
The defining feature of RRMHD compared to ideal relativistic magneto-hydrodynamics (MHD) is a finite conductivity $\sigma$, which is the inverse of the resistivity. In ideal relativistic MHD models of QGP, the conductivity is taken to be infinite, in other words, QGP is treated as a perfect conductor. However, Lattice QCD and pQCD calculations suggest the conductivity should be $\sigma=\mathcal{O}(0.1\sim 10)$ for the highest energy regions produced at RHIC and the LHC [7]. Additionally, a finite conductivity reduces the lifetime of the EM fields compared to ideal MHD and can introduce new phenomena like magnetic reconnection. Modeling QGP with a finite conductivity leads to interesting physics beyond the ideal MHD scenario.
@@ III Time evolution of the magnetic field
The magnetic field plays a major role in the charge-dependent directed flow $\Delta v_{1}$ at symmetric collisions. One role the magnetic field plays is to induce Faraday and Lorentz currents in the QGP. During the time evolution, the interplay of Lorentz + Coulomb + Faraday currents within the plasma leads to local imbalances in the electric chemical potential. Then that finite chemical potential becomes frozen onto the freezeout hypersurface. The local imbalances on the freezeout hypersurface caused by the time evolution of the magnetic field changes electric charge dependent observables like $\Delta v_{1}=v_{1}^{+}-v_{1}^{-}$. Before discussing the changes to $\Delta v_{1}$, we will focus on the time evolution of the magnetic field in our RRMHD model.
@@
  • We assume the electric conductivity of QGP as a constant value over all spacetime. Strictly speaking, dimensional analysis alone suggests the electric conductivity is temperature-dependent, which means it should change as the plasma expands and cools.
  • We ignore the pre-equilibrium dynamics that can create an initial electric charge distribution and change the configuration of the initial EM fields [26, 27]. Instead, we have chosen to initialize the EM fields using the solutions for the in-medium Maxwell equations [25]. Those solutions have been shown to overestimate the EM field values [24, 26, 27]. We have roughly estimated how that overestimation affects our model in Appendix C.
  • Last, the initial EM fields are created by the spectators of the collision, but during the time evolution we only include the plasma charges as a source of the EM fields. In other words, our model focuses on the behavior of charges inside the plasma and how they alone affect the time evolution of the EM fields. This may lead to an underestimation of the late time behavior of the EM fields, which we discuss near the end of Sec. V.
For a more complete description, we plan to resolve the first and third items in a future version of our model. The second item depends on an understanding of the pre-equilibrium dynamics, which has recently been an active area of research (e.g.,  [28, 29]). Some possible directions include color glass condensate (CGC), IP-Glasma, flux tube, and kinetic models.
@@
  • For $\sigma\geq 1$ the magnetic field is always over damped because the proper time is always greater than zero, $\tau>0$, in Milne coordinates.
  • For $0<\sigma<1$ the nature of the magnetic field damping depends on time. Starting in the region $\tau<1/(1-\sigma)$ the magnetic field is under-damped, then a smooth transition occurs to an over-damped region $\tau>1/(1-\sigma)$. The exact time of the transition depends on the value of $\sigma$.
This will be illustrated in Fig. 1 of Sec. V.
@@ IV Numerical setup and methods
For relativistic heavy-ion collisions, the symmetry of the system is best captured with Milne coordinates $(\tau,x,y,\eta)$, which have the metric $g_{\mu\nu}=\text{diag}(-1,1,1,\tau^{2})$. Accordingly, our RRMHD model solves Eqs.(1)-(5) using those coordinates with the volume factor $\sqrt{-g}=\tau$. We have adopted the optical Glauber model for the initial energy density profile and included an initial longitudinal tilt [31]. Even though we initialize the model after QGP is created at $\tau_{0}=0.4$ fm/$c$, the source of the initial EM fields is dominated by the collisions spectators [24]. Therefore, we calculate the initial EM fields using the analytic solution from [25], which is a solution of Maxwell’s equations in a medium with the ion charges as the source. The calculation is summarized in Appendix A. Overall, these are the same initial conditions as the previous works [23, 21, 20] and are listed in table 1.
@@
Numerically, we use a constant grid size of $168\times 168\times 72$ with $16.8\text{ fm}<(x,y)<16.8\text{ fm}$ and $7.2<\eta<7.2$, except for calculating the initial EM fields. Instead, for the initial EM fields, we adopt the following approach based on the analytic solution [25], which is summarized in Appendix A.
@@
    Assume the nuclei are very thin in the beam direction $z$ because of Lorentz contraction, such that their charge is distributed only in the transverse plane $x,y$. Then we can fix the $z$ position of the charges based on the collision kinematics.
@@
    At a grid point $\begin{pmatrix}x,&y,&\eta\end{pmatrix}$, calculate the EM fields by integrating over the contribution from all possible sources. Practically, this means we create a new grid for the EM sources in the transverse plane of size $200\times 200$ for $15.0\text{ fm}<(x,y)<15.0\text{ fm}$ and numerically integrate over it. This size ensures the area of both nuclei are integrated over.
The second step is repeated for all the grid points $\begin{pmatrix}x,&y,&\eta\end{pmatrix}$. For the numerical time evolution, we have adopted $C_{\text{CFL}}=0.05$ for the Courant-Friedrichs-Lewy (CFL) value.
Once the time evolution is completed, we calculate the final state hadron distribution using the Cooper-Frye formula [33].
$\frac{dN}{dYp_{T}dp_{T}d\phi}=\int_{\Sigma_{f}}f(x,p)p^{\mu}d\Sigma_{\mu}.$ (19)
We are ignoring any final state interactions that would be included in an afterburner because our current goal is to focus on the effects of the QGP conductivity. Our implementation of the Cooper-Frye formula constructs a 4-D hypersurface at the energy density boundary $\epsilon_{f}=0.15\text{ GeV/fm}^{3}$ [34]. We construct the local distribution $f(x,p)$ of Eq.(19) from the fluid velocities $v_{i}$, and charge density $q$ that is perpendicular to the hypersurface $d\Sigma(t,x,y,z)$,
where $Q$ is the quantized electric charge of the hadron.
For this work, the time evolution of the fluid is solved using an ideal-gas equation of state (EoS) 222 We have assumed the chemical potential $\mu_{Q}(x)$ is much smaller than the temperature of the fluid. Our assumption was based on the lack of baryon stopping at high-energy collisions observed near mid-rapidity and relating $\mu_{Q}=\mu_{B}/3$ [35, 36]. However, to mimic the experimentally observed splitting between the direct flow of hadrons, our model needs to include a finite chemical potential on the freezeout hypersurface. Then, to construct the potential on the hypersurface we need to introduce a new EoS that depends on $\mu_{Q}$. . However, on the freezeout hypersurface the distribution of Eq. (20) requires knowledge of the electric charge chemical potential $\mu_{Q}(x)$. To calculate $\mu_{Q}(x)$, we relate it to the local charge density $q(x)$ using a different EoS, i.e., Ref. [37], with the fluid’s local pressure $P$ and the thermodynamic relation $q=(\partial P/\partial\mu_{Q})$. Up to the linear approximation of $q(x)$ the relation is,
$\mu_{Q}(x)\simeq\frac{q(x)}{g_{h}T_{f}^{2}}$ (21)
where $g_{h}$ is the hadronic degrees of freedom. The final temperature in Eq. (20) is calculated for a massless boson gas,
$T_{f}=\left(\frac{30\epsilon_{f}(\hbar c)^{3}}{9\pi^{2}}\right)^{1/4}\approx 0.140\text{ GeV,}$ (22)
using the boson gas degrees of freedom. Finally, we integrate over the entire hypersurface to arrive at the final state hadron distribution.
@@ V RRMHD Results: Time evolution of the magnetic field
As discussed at the beginning of Sec. III, the time evolution of the magnetic field induces Lorentz + Faraday currents in QGP. Coulomb currents caused by the electric field of the positive spectators also appear, but for symmetric collisions largely cancel each other. Then the EM field component in the charge dependent directed flow $\Delta v_{1}$ for symmetric collisions is dominated by the Lorentz + Faraday currents Eq.(6), specifically the magnetic field. In particular, the $B_{y}$ component of the field has the largest contribution to $\Delta v_{1}$, which approximately appears as $\gamma B_{y}$ in the Faraday current and $\gamma u_{z}B_{y}$ in the Lorentz current [15]. Where $\gamma$ is the Lorentz factor of QGP in the local fluid rest frame and $u_{z}$ is the $z$ component of the fluid velocity.
In our RRMHD model, the time evolution of $B_{y}$ depends on the initial conditions and the electric conductivity $\sigma$ of QGP. As a starting point, we have used smooth initial solutions to Maxwell’s equations in a medium as discussed in Sec. IV that predicts a large initial field strength for $B_{y}$. That being said, solving the time evolution of the magnetohydrodynamics Eqs. (1)-(5) with different $\sigma$’s has an impact on $\Delta v_{1}$ through changes of $B_{y}$.
@@
Figure 1: Time evolution of the $y$-component of the magnetic field at the center of the grid using our RRMHD model. We have assumed the electric conductivity of the plasma $\sigma$ is a constant in spacetime, and the spectators of the collision only act as an initial source for the field. The magnetic field’s time evolution is similar to a non-linearly damped oscillator. For conductivity values between, $0<\sigma<1$ the magnetic field starts under damped and oscillates around zero. Larger values of $\sigma$ cause the magnetic field to be over damped, which means it never crosses zero.
The reduction in differences between the results for larger sigma values, like between $\sigma=10\text{ fm}^{-1}$ and $\sigma=100\text{ fm}^{-1}$, is partly due to numerical diffusion overwhelming the physical diffusion. This is a long-standing issue for resistive (relativistic) MHD models in astrophysics, where the conductivities can become relatively large, see Ref. [38] as an example of a detailed analysis. Here for each simulation, we keep the typical length-scale $L$ of the problem fixed and the order of the fluid velocities $u$ remains largely unchanged, but the conductivity increases by an order of magnitude. This means the magnetic Reynolds number $S\equiv\sigma L|u|$ also increases by an order of magnitude for each conductivity. When studying QGP, we limit the conductivity to be $\sigma<5\text{ fm}^{-1}$. This limit is justified by Lattice, pQCD, and AdS/CFT calculations and should keep the numerical diffusion smaller than the physical [4, 5, 6, 7, 8, 9].
A useful feature of the analytic solutions used for the initial EM fields, see Appendix  A and Sec. IV, is that it can be applied to all times. As a comparison in Fig. 2 the analytic solution for $B_{y}$ follows a similar time evolution as our RRMHD solution. The deviations after $\tau\sim 1.25$ fm/$c$ between the analytic and RRMHD solutions can be attributed to different EM sources used in each calculation. During the time evolution, our RRMHD model only considers QGP as a possible EM source $q$ in Eq.(5). Whereas, the analytic calculation includes all charges from the colliding ions, see Eq.(27) and the discussion after it. The difference between the two results may indicate our RRMHD model underestimates the EM fields, although the analytic solution is known to be an overestimation [24, 26]. At later times, we expect the charges of the spectators to have a limited effect because they have moved away from the collision region.
@@ VI RRMHD Results: Charge dependent directed flow
$v^{\pm}_{1}(Y)=\frac{\int p_{T}dp_{T}d\phi\cos\phi\frac{dN^{\pm}}{dYp_{T}dp_{T}d\phi}}{\int p_{T}dp_{T}d\phi\frac{dN^{\pm}}{dYp_{T}dp_{T}d\phi}},$ (24)
which is dependent on $Y$ the momentum-space rapidity and the hadron distribution is calculated using the Cooper-Frye formula of Eq.(19[33, 34]. Finally, we numerically integrate over the transverse momentum $p_{T}$ and angle $\phi$ in Eq.(24).
For our model, the source of non-vanishing $\Delta v_{1}$ is a nonzero electric chemical potential $\mu_{Q}$ on the freezeout hypersurface, which modifies the local thermal distribution $f(x,p)$ of Eq.(20). A nonzero chemical potential appears because locally the electric charge $q$ of the plasma is nonzero. From Eq.(5) Ohm’s law $q=-J^{\mu}u_{\mu}$, changes to the comoving charge density depend on the Lorentz and Faraday currents of Eq.(6), which are proportional to the conductivity $\sigma$.
Starting from an initial charge density of zero, Fig. 3 is a slice of the freezeout hypersurface at mid-rapidity in the transverse plane. The contours are along constant proper times $\tau$ starting from the edge of the freezeout hypersurface. The left panel is for a conductivity of $\sigma=0.0294\text{ fm}^{-1}$ and the right panel is for a larger conductivity of $\sigma=0.294\text{ fm}^{-1}$. The differences in the panels illustrate that the magnitude of $\mu_{Q}$ depends on $\sigma$, larger conductivities result in larger magnitudes of $\mu_{Q}$. This means the observable $\Delta v_{1}$ should depend on the conductivity of QGP via $\mu_{Q}$ and Eq.(20) the local thermal distribution.
@@
Figure 3: Electric chemical potential $\mu_{Q}$ that is accumulated on the freezeout hypersurface for $T_{f}=0.140$ GeV and $0.2>\eta>0.0$. Contours are over constant proper times of $\tau>0.4$ fm/$c$. The inner contours denote later times. Because we assumed the initial proton charges were equally distributed inside the ion, the transverse plane has reflection symmetry about zero on the $x$-axis and $y$-axis. The difference in magnitude of $\mu_{Q}$ between the panels is attributed to the value of QGPs electric conductivity $\sigma$.
@@
In Fig. 3, $\mu_{Q}$ has a reflection symmetry about $x=0$ and $y=0$ for that slice of the freezeout hypersurface. In contrast, Fig. 4 illustrates an asymmetry in $\mu_{Q}$ across the longitudinal axis $\eta=0$ on the freezeout hypersurface. This asymmetry is created in peripheral collisions, i.e., $b\neq 0$ fm, because the initial EM fields of the spectators break the reflection symmetry across $\eta=0$ [23]. Consequently, the charged directed flow of Eq.(24) is different for positively and negatively charged final state hadrons.
Figure 4: Longitudinal dependence of the electric chemical potential $\mu_{Q}$ that is accumulated on the freezeout hypersurface at $T_{f}=0.140$ GeV and $0.2>y>0.0$. The contours are equivalent to the ones in Fig.3, they are over constant proper times of $\tau>0.4$ fm/$c$. For peripheral collisions, the magnetic fields in the longitudinal plane $\eta$ are asymmetric about zero [22]. That leads to an asymmetry in $\mu_{Q}$ across $\eta$, which appears on the freezeout hypersurface. A 180 degrees rotation symmetry about the center $(0,0)$ exists because we assumed the initial proton changes were equally distributed inside the ions.
@@
An observable difference between the flow of protons and anti-protons is illustrated in Fig. 5, where the conductivity of QGP is varied across the panels. The rapidity $Y$ dependence of (anti-)proton charge dependent directed flow Eq.(24) separates in panels b and c. That separation is captured by a nonzero difference $\Delta v_{1}\neq 0$ from Eq.(23). This is a similar result as the observations at the STAR experiment [16].
@@
The negative linear slope of $\Delta v_{1}$ in panels b and c of Fig. 5 is a sign that the Faraday current has a greater net effect than the Lorentz current of Eq.(6). That result is in agreement with Ref. [15]. This shows the approximation of the Faraday current as $\gamma B_{y}$ and the Lorentz current as $\gamma u_{z}B_{y}$ at the beginning of Sec. III holds for our RRMHD model. Moreover, an order increase in the electric conductivity $\sigma$ brought a similar increase in the slope, $\Delta(dv_{1}/dy)$ suggesting this observable could be used to understand $\sigma$, the electric conductivity of QGP. Although, a more refined model is required to draw any quantitative conclusions.